On a certain winter day in Utah, the average atmospheric pressure is 738 torr. What is the molar density (in mol/L) of the air if the temperature is -18 degrees Celsius?
Answer #1
T = 273- 18 = 255 K
p = 738 / 760 = 0.971 atm
p / RT = molar density
0.971 / 0.0821 x 255 = 0.0464 mol/L